# Bi-Factorial Preference Optimization: Balancing Safety-Helpfulness in Language Models

### v1

@@ 4.4 Ablations
We validate the technical design of our algorithm in Table 4, showing that the shift parameter $\alpha$ and buffered training are effective in improving harmlessness.
In the BFPO $\alpha=0$ experiment, we set the shift parameter $\alpha$ to 0. The results indicate that, as illustrated in Section 3.4, the shift parameter $\alpha$ is useful in distinguishing unsafe data, and thus improves performance on generative tasks in harmlessness slightly. In the BFPO - w/o buffer experiment, we do not balance examples from the safety dataset and the helpful dataset, but instead mix the two datasets and randomly sample data from them. The lower harmlessness performance provides the evidence that buffered training helps mitigate the tension between helpfulness and harmlessness. Full results are provided in Section C.3.
@@ 5 Related Work
Alignment with Diverse Preferences. Traditional language model alignment methods [56, 22, 37] typically use a single reward or unified preference model.However, recent work suggests that human preferences are diverse and cannot be adequately represented by a single reward model. To address this, Chakraborty et al. [26] propose learning a mixture distribution for the reward using the EM algorithm, which they then apply in their MaxMin RLHF approach. Other research [52, 51, 27] explores training multi-objective reward models for the alignment stage. These methods primarily focus on improving reward models for RL based alignment. The most closely related work of supervised alignment methods is by Zhou et al. [21], who, despite advocating for direct policy optimization, still rely on training reward models. In contrast, our approach completely eliminates the two-stage training process and directly integrates multiple preferences into the supervised optimization.
Safety Alignment. Aligning large language models (LLMs) with both helpfulness and harmlessness is a specific case of addressing diverse preferences. To enhance safety, many researchers conduct additional safety data annotation alongside algorithm design. Llama2 [14] utilizes substantial amounts of human-labeled safety data and combines safety and helpfulness rewards by utilizing the safety reward as a threshold function for the helpfulness reward. [16, 57] engage in red teaming to gather extensive safety data and frame safety alignment as a conditioned Markov Decision Process (MDP) problem. Mu et al. [17] propose a rule-based reward as a complement for the common reward to improve the safety, which, although data-efficient, still requires human annotation and reward learning. In contrast, our method is fully automated and efficient, eliminating the need for human intervention in the safety alignment process. On the other hand, [58] propose generation-aware alignment, which improves the safety over different generation configurations. With our focus on improving safety under specific configurations, this work can be a strong complement to ours.
Safety Evaluation. Supervised benchmarks, such as OpenLLM [32] and BigBench [31], include datasets related to various aspects of safety, such as toxicity, truthfulness, morality, and social bias. Adversarial attack research [12] and red teaming efforts [57, 43] provide valuable toxic prompts to assess if models can generate harmless content in response to these prompts. To identify if the output content contains harmful information, some studies [13, 14] rely on human annotators, while others employ AI models like GPT-4 [59]. [43] offer fine-tuned Llama2 models to as harmful content classifier, offering an efficient alternative to GPT-4 in model-based evaluation.
@@ 6 Limitations and Discussion
ACKNOWLEDGEMENTS
This work is supported by the KAUST Competitive Research Grant (URF/1/4648-01-01), KAUST Baseline Fund (BAS/1/1685-01-01), and a UKRI grant Turing AI Fellowship (EP/W002981/1). AB has received funding from the Amazon Research Awards. We also thank the Royal Academy of Engineering and FiveAI.
@@ Appendix A Algorithm
Algorithm 1 shows the BFPO algorithm. As mentioned in Section 2, in practice, we refer to datasets related to safety topics, collected through red teaming, as safety datasets. A typical safety dataset will contain a safety label $I_{\text{safe}}(y)$, which is the binary label indicating whether the response $y$ is harmful, as well as the preference label $I_{\text{help}}(y\succ y^{\prime})$ in terms of helpfulness. If a certain safety dataset does not provide helpfulness labels, we can use the ranking models, like PairRM [55], as discussed in Section 4.3, to generate the pairwise helpfulness labels. We refer to datasets designed to improve the helpfulness of the model as helpfulness datasets. A typical helpfulness dataset will contain the helpfulness preference labels $I_{\text{help}}(y\succ y^{\prime})$. Since most helpfulness data undergoes human filtering, the responses are usually safe. Therefore, we assign the safety label $I_{\text{safe}}(y)=1$ to all responses in the helpfulness dataset.
@@
8:         Compute $h(h^{hw},h^{hl})$ with Equation 5
@@ B.1 Notation
To establish optimal solution, we follow Azar et al. [19] to leverage the following lemma.
Lemma B.2 (Rafailov et al. [18], Azar et al. [19]).
@@
$\text{argmax}_{\delta\in\Delta(\mathcal{S})}\mathcal{L}_{\tau}(\delta)$
@@
Lemma B.3 (Theorem 2 in Azar et al. [19]).
@@
Now we prove that $\pi_{s^{*}}$ is the unique global minima to $\mathcal{L}(s)$. Since $\pi_{s}$ is a surjective continuous mapping from $s$ to $\pi$, then every local minima $\pi$ to $\mathcal{L}(\pi)$ corresponds to a set of $s$ that minimizes $\mathcal{L}(s)$. The uniquess of $s^{*}$ will deduce that $\pi^{*}$ is the unique optimal solution to Equation 18 and concludes the proof. Consider $s^{\prime}=s^{*}+r\cdot\Delta s$, where the only $r$ is the radius and $\Delta s$ is the direction under the polar coordinate. The only direction that not increase $\mathcal{L}(s^{\prime})$ away from 00 is $e=(\frac{1}{n},\dots,\frac{1}{n})$ ([60], Chap. 3). However, we have
@@
Now we provide the proof of Theorem 3.1, most of which follows Azar et al. [19].
@@
$\small\begin{split}&g_{I}(y,y^{\prime})=B_{3}\Big{(}B_{1}\big{(}I_{\text{safe}}(y)I_{\text{help}}(y\succ y^{\prime})+I_{\text{safe}}(y^{\prime})I_{\text{help}}(y^{\prime}\succ y)\big{)}\\ &-\big{(}I_{\text{safe}}(y)I_{\text{help}}(y^{\prime}\succ y)+I_{\text{safe}}(y^{\prime})I_{\text{help}}(y\succ y^{\prime})\big{)}+B_{2}\Big{)}\cdot\Big{(}2I_{\text{help}}(y\succ y^{\prime})-1\Big{)},\end{split}$ (21)
@@ B.5 Discussion of the Property of $g_{I}$
This means that for the same data point, regardless of the order of $y$ and $y^{\prime}$, we are always driving $h_{\pi}(y,y^{\prime})$ to the same value. In contrast, in IPO [19], different orders will push $h_{\pi}(y,y^{\prime})$ to different values, i.e., they form two different optimization problems:
@@ C.2 Details of Baselines
  • Llama2 [14] trains the safety reward $r_{\text{safe}}$ and the helpfulness reward $r_{\text{help}}$ separately, and defines the global reward $g$ as a combination of these rewards, i.e.,
@@
  • Beaver [16] trains the safety reward $r_{\text{safe}}$ and the helpfulness reward $r_{\text{help}}$ separately, and defines the final RLHF objective as the dual optimization problem of the conditional RLHF, obtained by Lagrangian dual transformation, i.e.,
@@
  • RBR [17] requires separate reward models, $r_{\phi_{1}},\dots,r_{\phi_{k}}$, for each objective, and propose to learn the weight for each objective, i.e.,
@@
  • SteerLM [52] trains models to generate response according to a specific reward vector $r=(r_{1},r_{2},r_{3},\dots,r_{k})$. They first train a model to predict the score for each objective in a dataset. Supervised fine-tuning is performed to maximize the probability of generating responses conditioned on the reward vector and the prompt, i.e.,
@@
  • MORL [51] trains reward models for each objective separately, and defines the global reward $g$ as a combination of rewards, i.e.,
@@
  • ArmoRM [27] applies the same training strategy as MORL, but uses a single publicly available reward model, ArmoRM-Llama3-8B-v0.1 [54], to provide the reward scores for all objectives.
  • MODPO [21] trains margin reward models $r_{i},i=1,\dots,k$ for each objective separately, and performs supervised fine-tuning with the objective,
@@
  • MinMaxRLHF [26] addresses the scenario where different annotators $h$ may have preferences for different objectives. The algorithm uses the EM algorithm to learn the distribution of rewards for multiple objectives. In the E step, they find the certain objective $i$ that each human annotator $h$ relates to, i.e.,
    $\mathcal{I}_{h}=\text{argmax}_{i}\Pi_{x,y,y^{\prime},h}\frac{\exp(r_{\phi_{i}}(x,y))}{\exp(r_{\phi_{i}}(x,y))+\exp(r_{\phi_{i}}(x,y^{\prime}))},$
@@
Among these methods, MODPO is highly inefficient since it requires separate RLHF for each objective. Other methods typically use a linear combination of reward scores for multiple objectives or one reward as a threshold for others. For the combination of thresholding, the global function can be approximated by the multiplication of rewards for each objective when the reward scores are on the same scale. Maximizing the multiplication of rewards has the same effect as maximizing the minimum reward. Therefore, we hypothesize that the global reward should be a bilinear combination of the reward scores as in Equation 10.
C.3 Full Experiment Results
Table 6 shows the full results of the open-sourced models, our baselines, and our models on the benchmarks. Here are the details of each open-sourced models:

### v2

@@ 4.4 Ablations
We validate the technical design, especially the shifting parameter $\alpha$ and the buffer training in Table 4. In the BFPO $\alpha=0$ experiment, we set the shift parameter $\alpha$ to 0. The results indicate that, as illustrated in Section 3.4, the shift parameter $\alpha$ is useful in distinguishing unsafe data, and thus improves performance on generative tasks in harmlessness slightly. In the BFPO - w/o buffer experiment, we do not balance examples from the safety dataset and the helpful dataset, but instead mix the two datasets and randomly sample data from them. The lower harmlessness performance provides the evidence that buffered training helps mitigate the tension between helpfulness and harmlessness. Full results and detailed ablation are provided in Section C.4 and Section C.3.
Other hyper-parameters, like $\tau,B_{1}$, are set based on either our theoretical understanding or the past work (Tang et al., 2024), and the fine-tuning strategy is orthogonal to the algorithm, while we further include their ablation in Section C.3.
@@ 5 Related Work
Alignment with Diverse Preferences. Traditional language model alignment methods (Christiano et al., 2017; Stiennon et al., 2020; Hendrycks et al., 2021) typically use a single reward or unified preference model.However, recent work suggests that human preferences are diverse and cannot be adequately represented by a single reward model. To address this,  Chakraborty et al. (2024) propose learning a mixture distribution for the reward using the EM algorithm, which they then apply in their MaxMin RLHF approach.  Dong et al. (2023); Ramé et al. (2023); Wang et al. (2024b) explore training multi-objective reward models for the alignment stage. These methods primarily focus on improving reward models for RL based alignment. The most closely related work of supervised alignment methods is by  Zhou et al. (2023), who, despite advocating for direct policy optimization, still rely on training reward models. In contrast, our approach completely eliminates the two-stage training process and directly integrates multiple preferences into the supervised optimization.
Safety Alignment. Aligning large language models (LLMs) with both helpfulness and harmlessness is a specific case of addressing diverse preferences. To enhance safety, many researchers conduct additional safety data annotation alongside algorithm design.  Touvron et al. (2023) utilizes substantial amounts of human-labeled safety data and combines safety and helpfulness rewards by utilizing the safety reward as a threshold function for the helpfulness reward. Dai et al. (2024); Ji et al. (2024) engage in red teaming to gather extensive safety data and frame safety alignment as a conditioned Markov Decision Process (MDP) problem.  Mu et al. (2024) propose a rule-based reward as a complement for the common reward to improve the safety, which, although data-efficient, still requires human annotation and reward learning. In contrast, our method is fully automated and efficient, eliminating the need for human intervention in the safety alignment process. On the other hand, Huang et al. (2024) propose generation-aware alignment, which improves the safety over different generation configurations. With our focus on improving safety under specific configurations, this work can be a strong complement to ours.
Safety Evaluation. Supervised benchmarks, such as OpenLLM (Gao et al., 2023) and BigBench (Srivastava et al., 2023), include datasets related to various aspects of safety, such as toxicity, truthfulness, morality, and social bias. Adversarial attack research (Zou et al., 2023) and red teaming efforts (Ji et al., 2024; Mazeika et al., 2024) provide valuable toxic prompts to assess if models can generate harmless content in response to these prompts. To identify if the output content contains harmful information, some studies (Bai et al., 2022; Touvron et al., 2023) rely on human annotators, while others employ AI models like GPT-4 (Wang et al., 2024a). Mazeika et al. (2024) offer fine-tuned Llama2 models to as harmful content classifier, offering an efficient alternative to GPT-4 in model-based evaluation.
@@ 6 Limitations and Discussion
Aknowledgement
This work is supported by a KAUST CRG (URF/1/4648-01-01) and a UKRI grant Turing AI Fellowship (EP/W002981/1). A. Bibi acknowledges the Google Gemma 2 Academic Award 2024. We also thank the Royal Academy of Engineering.
@@ Appendix A Algorithm
Algorithm 1 shows the BFPO algorithm. As mentioned in Section 2, in practice, we refer to datasets related to safety topics, collected through red teaming, as safety datasets. A typical safety dataset will contain a safety label $I_{\text{safe}}(y)$, which is the binary label indicating whether the response $y$ is harmful, as well as the preference label $I_{\text{help}}(y\succ y^{\prime})$ in terms of helpfulness. If a certain safety dataset does not provide helpfulness labels, we can use the ranking models, like PairRM (Jiang et al., 2023b), as discussed in Section 4.3, to generate the pairwise helpfulness labels. We refer to datasets designed to improve the helpfulness of the model as helpfulness datasets. A typical helpfulness dataset will contain the helpfulness preference labels $I_{\text{help}}(y\succ y^{\prime})$. Since most helpfulness data undergoes human filtering, the responses are usually safe. Therefore, we assign the safety label $I_{\text{safe}}(y)=1$ to all responses in the helpfulness dataset.
@@
8:         Compute $h(y^{hw},y^{hl})$ with Equation 5
@@ B.1 Notation
To establish optimal solution, we follow Azar et al. (2024) to leverage the following lemma.
Lemma B.2 Rafailov et al. (2023),  Azar et al. (2024)).
@@
$\operatorname*{arg\,max}_{\delta\in\Delta(\mathcal{S})}\mathcal{L}_{\tau}(\delta)$
@@
Lemma B.3 (Theorem 2 in  Azar et al. (2024)).
@@
Now we prove that $\pi_{s^{*}}$ is the unique global minima to $\mathcal{L}(s)$. Since $\pi_{s}$ is a surjective continuous mapping from $s$ to $\pi$, then every local minima $\pi$ to $\mathcal{L}(\pi)$ corresponds to a set of $s$ that minimizes $\mathcal{L}(s)$. The uniquess of $s^{*}$ will deduce that $\pi^{*}$ is the unique optimal solution to Equation 18 and concludes the proof. Consider $s^{\prime}=s^{*}+r\cdot\Delta s$, where the only $r$ is the radius and $\Delta s$ is the direction under the polar coordinate. The only direction that not increase $\mathcal{L}(s^{\prime})$ away from 00 is $e=(\frac{1}{n},\dots,\frac{1}{n})$ (Boyd & Vandenberghe (2004), Chap. 3). However, we have
@@
Now we provide the proof of Theorem 3.1, most of which follows  Azar et al. (2024).
@@
$\begin{split}&g_{I}(y,y^{\prime})=B_{3}\Big{(}B_{1}\big{(}I_{\text{safe}}(y)I_{\text{help}}(y\succ y^{\prime})+I_{\text{safe}}(y^{\prime})I_{\text{help}}(y^{\prime}\succ y)\big{)}\\ &-\big{(}I_{\text{safe}}(y)I_{\text{help}}(y^{\prime}\succ y)+I_{\text{safe}}(y^{\prime})I_{\text{help}}(y\succ y^{\prime})\big{)}+B_{2}\Big{)}\cdot\Big{(}2I_{\text{help}}(y\succ y^{\prime})-1\Big{)},\end{split}$ (21)
@@ B.5 Discussion of the Property of $g_{I}$
This means that for the same data point, regardless of the order of $y$ and $y^{\prime}$, we are always driving $h_{\pi}(y,y^{\prime})$ to the same value. In contrast, in IPO (Azar et al., 2024), different orders will push $h_{\pi}(y,y^{\prime})$ to different values, i.e., they form two different optimization problems:
@@ C.2 Details of Baselines
  • Llama2 (Touvron et al., 2023) trains the safety reward $r_{\text{safe}}$ and the helpfulness reward $r_{\text{help}}$ separately, and defines the global reward $g$ as a combination of these rewards, i.e.,
@@
  • Beaver (Dai et al., 2024) trains the safety reward $r_{\text{safe}}$ and the helpfulness reward $r_{\text{help}}$ separately, and defines the final RLHF objective as the dual optimization problem of the conditional RLHF, obtained by Lagrangian dual transformation, i.e.,
@@
  • RBR (Mu et al., 2024) requires separate reward models, $r_{\phi_{1}},\dots,r_{\phi_{k}}$, for each objective, and propose to learn the weight for each objective, i.e.,
@@
  • SteerLM (Dong et al., 2023) trains models to generate response according to a specific reward vector $r=(r_{1},r_{2},r_{3},\dots,r_{k})$. They first train a model to predict the score for each objective in a dataset. Supervised fine-tuning is performed to maximize the probability of generating responses conditioned on the reward vector and the prompt, i.e.,
@@
  • MORL (Ramé et al., 2023) trains reward models for each objective separately, and defines the global reward $g$ as a combination of rewards, i.e.,
@@
  • ArmoRM (Wang et al., 2024b) applies the same training strategy as MORL, but uses a single publicly available reward model, ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024c), to provide the reward scores for all objectives.
  • MODPO (Zhou et al., 2023) trains margin reward models $r_{i},i=1,\dots,k$ for each objective separately, and performs supervised fine-tuning with the objective,
@@
  • MinMaxRLHF (Chakraborty et al., 2024) addresses the scenario where different annotators $h$ may have preferences for different objectives. The algorithm uses the EM algorithm to learn the distribution of rewards for multiple objectives. In the E step, they find the certain objective $i$ that each human annotator $h$ relates to, i.e.,
    $\mathcal{I}_{h}=\operatorname*{arg\,max}_{i}\Pi_{x,y,y^{\prime},h}\frac{\exp(r_{\phi_{i}}(x,y))}{\exp(r_{\phi_{i}}(x,y))+\exp(r_{\phi_{i}}(x,y^{\prime}))},$
@@
Among these methods, MODPO is highly inefficient since it requires separate RLHF for each objective. Other methods typically use a linear combination of reward scores for multiple objectives or one reward as a threshold for others. For the combination of thresholding, the global function can be approximated by the multiplication of rewards for each objective when the reward scores are on the same scale. Maximizing the multiplication of rewards has the same effect as maximizing the minimum reward. Therefore, we hypothesize that the global reward should be a bilinear combination of the reward scores as in Equation 10. As such, in the experiment section, we select MORL as a representative for this line of approach.
C.3 Additional Ablation Studies
We include additional ablation studies of other hyper-parameters in our algorithms. Full results are in Table 10.
LoRA finetuning.
We follow the setting of Section 4.2 and conduct additional experiments to compare our method and the best performed baseline, IPO, with LoRA fine-tuning. During the training, we apply the same training hyper-parameters to both algorithms, like learning rate, training epochs, beta, and so on. The results are in Table 6. Results show that BFPO consistently outperformed the baselines when training with LoRA. However, we observed that LoRA training required additional hyperparameter tuning, which posed challenges due to the limited time. Consequently, both methods achieved lower overall performance and worse balance compared to selective fine-tuning.
LoRA Fine-tuning Selective Fine-tuning
Helpfulness Harmlessness Helpfulness Harmlessness
Alpaca Disc. Gen. Savg. Alpaca Disc. Gen. Savg.
IPO 6.14 58.05 93.97 76.1 13.15 58.41 89.76 74.09
BFPO 7.77 64.36 94.73 79.54 3.33 59.09 95.24 77.16
Table 6: Comparison of Helpfulness and Harmlessness Metrics with LoRA finetuning
Table 7: Ablation Study of $\alpha$ in Equation 15
$\alpha$ Helpfulness Harmlessness
Alpaca Disc. Gen. Savg.
0.1 13.61 59.81 87.39 73.60
0.3 14.06 60.31 91.73 76.02
0.5 13.33 59.09 95.24 77.16
0.7 9.01 57.34 96.28 76.81
0.9 7.21 56.44 96.66 76.55
Table 8: Ablation Study of $\tau$ in Equation 15
$\tau$ Helpfulness Harmlessness
Alpaca Disc. Gen. Savg.
0.01 13.33 59.09 95.24 77.16
0.1 6.4 55.44 81.45 68.44
0.5 6.53 54.01 78.14 66.07
1.0 6.74 54.10 77.52 65.81
The hyperparameter $\alpha$.
The hyperparameter controls the label values (represent the difference of the preference of a pair of response) of the four cases in Figure 3. To ensure the desired behavior, that helpful-safe responses are preferred over helpless-safe ones (case 2 in Figure 3 yields a positive value) and that helpful-unsafe responses are not preferred over helpless-unsafe ones (Case 3 in Figure 3 yields a negative value), we constrain $\alpha\in(0,1)$. When $\alpha=0.5$, the label values for the four cases are $1,0.5,-0.5,-1$, where the absolute label values are symmetric for positive and negative pairs. As $\alpha$ increases, the absolute label values in case 1,2 in Figure 3 decrease, and the absolute label values in case 3,4 in Figure 3 increase. In other words, positive pairs will have smaller differences and negative pairs will have larger differences.
In the ablation study, we follow the experiment of Section 4.2 with values of $0.1,0.3,00.5,0.7,0.9$ to systematically explore its effects. The results in Table 8 show that higher $\alpha$ values reduce distinctions between positive pairs, particularly helpful-safe vs. non-helpful-safe, leading to a lower helpfulness score. However, it increases distinctions between negative pairs, especially helpful-unsafe vs. non-helpful-safe, resulting in improved harmlessness, particularly in generative tasks.
The hyperparameter $\tau$.
The hyperparameter $\tau$ is the coefficient of the KL term in Equation 7, which prevents the policy from deviating from the reference policy. In practice, it is important to note that is more related to the training and convergence (as shown in Figure 5) rather than being a core component to balance the safety and helpfulness.
In our experiments, we follow the settings from Tunstall et al. (2023), where $\tau=0.01$ is used. This value is applied consistently across all baselines to ensure a fair comparison. For the ablation study, we adopt values inspired by Tang et al. (2024), specifically $\tau=0.01,0.1,0.5,1.0$. The results in Table 8 indicate that performance can vary significantly with different $\tau$ values. With different $\tau$, other training hyperparameters, like learning rate, training iterations also need to be carefully chosen.
Ablation on $B_{1},B_{2},B_{3}$.
both $B_{1},B_{3}$ must be positive. For this ablation study, we explore the following values $B_{3}=2,1,\frac{1}{2},\frac{1}{4}$. Given the constraint $B_{3}(B_{1}-1)=1$ as in Equation 33, the corresponding values of $B_{1}$ are determined for each $B_{3}$. Additionally, $B_{2}$ is adjusted to balance the cases described in Figure 3 (Case 2 and Case 3). The experiment results are in Table 9.
When $B_{3}$ is smaller, the label differences for cases 1,2 and 3,4 in Figure 3 become less pronounced. For example, in Cases 1 and 2, the pairs (helpful-safe, non-helpful-unsafe) and (helpful-safe, non-helpful-safe) have smaller differences in their label values. This means there is less distinction in whether the non-helpful response is safe or not. As a result, the model shows slightly worse performance in helpfulness but performs better in safety. When $B_{3}$ is larger, the label differences for the aforementioned two cases become more distinct, and the label value for (helpful-safe, non-helpful-unsafe) becomes significantly higher. This leads the model to prioritize safety more strongly, which results in improved safety performance but a sacrifice in helpfulness.
To conclude, larger $B_{3}$ values emphasize safety at the expense of helpfulness, while proper values allow for more balanced performance across both objectives.
Table 9: Ablation study for $B_{1},B_{2},B_{3}$ in Equation 9
$B_{3}$ $B_{1}$ $B_{2}$ Values for four cases in Figure 3 Helpfulness Harmlessness
Alpaca Disc. Gen. Savg.
2 1.5 -0.25 2.5,0.5,-0.5,-0.25 9.00 58.67 95.47 77.07
1 2 -0.5 1.5,0.5,-0.5,-1.5 11.36 60.28 95.12 77.70
0.5 3 -1 1,0.5,-0.5,-1 13.33 59.09 95.24 77.16
0.25 5 -2 0.75, 0.5, -0.5, -0.75 13.15 59.63 94.25 76.94
C.4 Full Experiment Results